OpenStudy (anonymous):

What are the factors of 25x2 + 40x + 16? (5x + 4)(5x + 4) (5x − 4)(5x − 4) (5x + 2)(5x + 8) (5x − 8)(5x − 2)

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

A?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Perfect square trinomial: a^2 + 2ab + b^2 = (a + b)(a + b) = (a + b)^2 25x^2 - 40x + 16 (5x - 4)(5x - 4) (5x - 4)^2 <===ANSWER

OpenStudy (nincompoop):

you don't sound so sure

OpenStudy (anonymous):

Awesome! Thanks guys! and @nincompoop I'm just getting the hang of this haha.

OpenStudy (anonymous):

@Gerardie it's +40x